SMOOTHLY PARAMETERISED CECH COHOMOLOGY OF COMPLEX MANIFOLDS

TOBY BAILEY MICHAEL EASTWOOD SIMON GINDIKIN

Abstract. A Stein covering of a complex manifold may be used to realise its analytic cohomology in accordance with the Cech theory. If, however, the Stein covering is parameterised by a smooth manifold rather than just a discrete set, then we construct a cohomology theory in which an exterior derivative replaces the usual combinatorial Cech differential. Our construction is motivated by integral geometry and the representation theory of Lie groups.